OpenStudy (anonymous):

8 * 63 = 8 * ____ + 8 * 3= ____

OpenStudy (chosenmatt):

welcome to openstudy

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

What are the answer choices??

OpenStudy (anonymous):

\[8*63=8*\left( 60+3 \right)=8*60+8*3\]

OpenStudy (anonymous):

the answers are 20 and 504 i got that by doing 8*20+8*3=504 and 8*63=504

OpenStudy (anonymous):

8*20=160 8*3=24 160+24=184

OpenStudy (anonymous):

8*63=504 8*60=480 8*3=24 480+24=504